# Learning-Augmented Robust Algorithmic Recourse

### v2

@@ 1 Introduction
Machine learning models are deployed even in sensitive domains such as lending or hiring, e.g., financial institutions use these models to determine whether someone should receive a loan. Given the major impact of these decisions on people’s lives, a plethora of work in responsible machine learning aims to make these models fair [11, 5, 32, 83], transparent [44, 65], and explainable [64, 49, 68]. A notable line of work along this direction, called algorithmic recourse [76, 71], provides each individual who received an undesirable label (e.g., one whose loan request was denied) with a minimum-cost improvement to achieve the desired label.
An important weakness of much of the work on algorithmic recourse is the assumption that models are fixed and do not change. In practice, models are periodically updated to reflect changes in data or environment, causing a recourse to become invalid, i.e., following it may not yield a desirable outcome [20]. To alleviate this problem, Upadhyay et al. [70] proposed a framework that is robust to adversarial model changes and provided an algorithm, ROAR, to compute robust recourses. However, guaranteeing robustness in such a setting leads to significantly increased cost [59].
Our main goal is to design more practical recourse algorithms that balance robustness and cost. To achieve this goal, we move beyond the adversarial setting and leverage the learning-augmented framework [53], which has been used in a surge of recent work to overcome the limitations of adversarial (i.e., worst-case) analysis. Specifically, rather than assuming that the designer has no information regarding the model changes, we assume they can at least formulate some predictions regarding these changes. However, these predictions are unreliable, so our goal is to compute recourses that are near-optimal if the predictions are accurate (consistency), while maintaining good performance even in the worst-case, i.e., even when the predictions are highly inaccurate (robustness).
Our Results We first adapt the learning-augmented framework and the objectives of consistency and robustness to algorithmic recourse, and we provide a computationally efficient algorithm for optimizing these two objectives, as well as their convex combinations. We then experimentally evaluate the performance of this algorithm on a variety of both linear and non-linear models. We also formally prove that our algorithm is optimal for both robustness and consistency for any generalized linear model. Note that this is a non-convex problem and, to the best of our knowledge, this is the first optimal algorithm for any robust recourse problem (prior work used algorithms which, as we show, return local rather than global optima). Therefore, this is a contribution to the broader robust recourse literature, beyond the learning-augmented setting.
@@ 1.1 Related Work
Other work on robust recourse following up on ROAR includes RBR [56], which provides robustness with respect to data shifts, and RoCourseNet, which simultaneously optimizes for the model and robust recourse (though the initial model is fixed in our approach). The learning-augmented framework is applied in domains such as algorithm and data structures design [50, 62, 43], mechanism design [2, 80], and privacy-preserving analysis [41]. See Appendix A.
@@ 2 Preliminaries
Consider a predictive model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\theta\in\Theta\subseteq\mathbb{R}^{d}$, which maps instances (e.g., loan applicants) from a feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$ to an outcome space $\mathcal{Y}=\{0,1\}$. The values of 0 and 1 represent undesirable and desirable outcomes (e.g., loan denial or approval), respectively. If a model $f_{\theta_{0}}$ yields an undesirable outcome for some instance $x_{0}\in\mathcal{X}$, i.e., $f_{\theta_{0}}(x_{0})=0$, the goal is to suggest an optimal recourse: the least costly way to modify $x_{0}$ (how an applicant should strengthen their application) so that the resulting instance $x^{\prime}$ would achieve the desirable outcome under $f_{\theta_{0}}$, i.e., $f_{\theta_{0}}(x^{\prime})=1$. Given a cost function $c:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ quantifying the cost $c\left(x^{\prime},x_{0}\right)$ required to modify $x_{0}$ to $x^{\prime}$, the recourse is defined as the following optimization problem [70]:
@@
where $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ is a convex loss function that is decreasing in its first argument (such as binary cross entropy or squared loss) and captures the extent to which the condition $f_{\theta_{0}}(x^{\prime})=1$ is violated and $\lambda\geq 0$ is a regularizer that balances the degree of violation from the desirable outcome and the cost of modifying $x_{0}$ to $x^{\prime}$. The regularizer $\lambda$ can be decreased gradually until the desired outcome is reached. In this work, following the approach of [70, 63], we assume the cost function is the $L^{1}$ distance, i.e. $c(x,x^{\prime})=\|x-x^{\prime}\|_{1}$.
@@
Equation (1) assumes that the parameters of the model remain the same over time, but in practice, predictive models may be periodically retrained and updated [70]. These updates can cause a recourse that is valid in the original model (i.e., one that would lead to the desirable outcome in that model) to become invalid in the updated model [21, 13]. It is, therefore, natural to require a recourse solution whose validity is robust to (slight) changes in the model parameters.
In line with prior work on robust recourse [70, 13], we assume that the parameters of the updated model can be any $\theta^{\prime}\in\Theta_{\alpha}$, where $\Theta_{\alpha}=\{\theta:\|\theta-\theta_{0}\|_{\infty}\leq\alpha\}\subseteq\Theta$ is a “neighborhood” around the parameters, $\theta_{0}$, of the original model, defined using the $L^{\infty}$ distance and a known parameter $\alpha$. Given $\theta_{0}$ and $\Theta_{\alpha}$, the robust solution would be to choose a recourse $x_{r}$ that minimizes the price assuming the parameters of the updated model $\theta^{\prime}\in\Theta_{\alpha}$ are chosen adversarially, i.e.,
@@ 3 Learning-Augmented Framework
Choosing the robust recourse $x_{r}$ according to (3) optimizes the price against an adversarially chosen $\theta^{\prime}\in\Theta_{\alpha}$, but this price may be much higher than the optimal price if we knew $\theta^{\prime}$. This is due to the overly pessimistic assumption that the designer has no information regarding the realized $\theta^{\prime}\in\Theta_{\alpha}$. To overcome similarly pessimistic results in other domains, the learning-augmented framework [53] provides a more refined analysis by assuming the designer is equipped with an unreliable prediction and then seeks to achieve near-optimal performance whenever the prediction is accurate while simultaneously maintaining some robustness even if the prediction is arbitrarily inaccurate.
@@
The robustness evaluates the worst-case price of $x^{\prime}$ against an adversarial change of the model and then compares it to the price of $x_{r}$. The robustness is always at least 0 (achieved by $x^{\prime}=x_{r}$), and lower is more desirable. While prior work [70] measures robustness in absolute terms, we evaluate it relative to $x_{r}$ to enable a direct comparison between robustness and consistency.
@@ 3.1 Computing Robust/Consistent Recourses
We next provide Algorithm 1 to solve Equation (7). We introduce a few additional notations. We use $\operatorname{sgn}(s)$ which is 1 when $s>0$, 0 when $s=0$, and 1-1 when $s<0$. When applied to a vector, the $\operatorname{sgn}$ is applied element-wise. For an integer $n\in\mathbb{N}$, $n]:=\{1,\ldots,n\}$.
@@
4:  Initialize worst-case $\theta^{\prime}[i]$ based on $x[i]$ and $\theta[i]$
5:  $\textsc{Active}\leftarrow\textsc{Active}\setminus\{i\}$ if $x^{\prime}_{i}$ cannot improve (7)
@@
7:  $i,\Delta\leftarrow$ FindOptimalDimensionAndUpdate($x_{0}$, $\theta$, $\ell$, $c$, $\alpha$, $\beta$, $\hat{\theta}$, $x^{\prime}$)
@@
11:   $x^{\prime}[i]\leftarrow x^{\prime}[i]+\Delta$ $\triangleright$ Update x’[i]
@@
Input: $x_{0}$, $\theta$, $\ell$, $c$, $\alpha$, $\beta$, $\hat{\theta}$, and $x^{\prime}$Output: $i,\Delta$
@@
4:  $x^{*}\in\arg\min_{x:x[j]=x^{\prime}[j]\forall j\neq i}K(x)$ $\triangleright$ Minimizer of $K$ only if dimension $i$ could change
5:  if $K(x^{\prime})-K(x^{*})>0$ and ($x^{\prime}[i]=x_{0}[i]$ or $\operatorname{sgn}(x^{\prime}[i]-x_{0}[i])=\operatorname{sgn}(x^{*}[i]-x^{\prime}[i]))$ then
6:   $O[i]\leftarrow K(x^{\prime})-K(x^{*})$ $\triangleright$ Only update if the change is consistent with prior changes in $x^{\prime}[i]$
@@
Algorithm 1 starts by approximating $f_{\theta}$ locally at $x_{0}$ with a linear model e.g., using LIME [64]. The algorithm then computes the worst-case model $\theta^{\prime}$ for the “default” recourse of $x^{\prime}=x_{0}$ (lines 3-5). See Appendix C. Then, facing $\theta^{\prime}$, the algorithm greedily modifies $x^{\prime}$ while simultaneously updating $\theta^{\prime}$ to ensure that it remains the worst-case model for $x^{\prime}$ (lines 6-17). More specifically, in each iteration of the while loop, for the current $\theta^{\prime}$, the algorithm first calls the subroutine FindOptimalDimensionAndUpdate (line 1), implemented in Algorithm 2. This subroutine searches over each dimension and returns the dimension $i$ and changes in that dimension $\Delta$ which minimizes the objective for a fixed $\theta^{\prime}$ and assuming all the other coordinates are fixed. After computing $\Delta$, Algorithm 1 terminates if $\Delta$ is 0, meaning no improvement can be done (line 9). If the change $\Delta$ does not cause $x^{\prime}[i]$ to flip its sign, the update is applied and the adversarial model $\theta_{0}$ does indeed remain fixed (lines 11). On the other hand, if the recommended change flips the sign of $x^{\prime}[i]$, this would cause the adversarial response to change as well. The algorithm instead applies this change all the way up to $x^{\prime}[i]=0$ and updates $\theta^{\prime}$ accordingly (lines 13-15). If, during this process, $x^{\prime}[i]$ becomes 0 and an update of the adversarial model $\theta^{\prime}[i]$ could cause its sign to flip, then no further change in this dimension is allowed i.e., $i$ is removed from the Active set (line 17).
Algorithm 1 is efficient since (1) the linear approximation of $f_{\theta_{0}}$ can be computed in polynomial time [64], and (2) the while loop runs for $O(d)$ iterations and the running time of each iteration is dominated by the run time of subroutine FindOptimalDimensionAndUpdate. This subroutine runs in polynomial time (e.g., by running at most $d$ one-dimensional gradient descents [14]). See Appendix C for more details, and Appendix D.7 for the scalability of our approach to a larger dataset.
@@
Note that $\beta=1$ corresponds to computing the optimal robust recourse. Even for generalized linear models, the objective function in Equation (7) is non-convex when $\beta=1$ (see Appendix C.1). Hence, prior gradient-based approaches [70, 56] cannot guarantee optimality. To the best of our knowledge, Algorithm 1 is the first optimal algorithm for any robust recourse formulation.
The idea of approximating a non-linear function locally has also been used in prior work [70, 63]. In Section 4, we empirically compare the performance of our algorithm to prior work in linear and non-linear settings. Moreover, there may be feasibility constraints on recourses, or the data may contain categorical features. To handle such cases, the recourse of Algorithm 1 can be post-processed (e.g., by projection) to guarantee feasibility [70, 56, 29]. See Appendix D.4.

### v3

@@ 1 Introduction
Machine learning models are deployed even in sensitive domains such as lending or hiring, e.g., financial institutions use these models to determine whether someone should receive a loan. Given the major impact of these decisions on people’s lives, a plethora of work in responsible machine learning aims to make these models fair (Berk et al., 2021; Barocas et al., 2019; Hardt et al., 2016; Zafar et al., 2017), transparent (Lakkaraju et al., 2016; Rudin, 2019), and explainable (Ribeiro et al., 2016; Lundberg and Lee, 2017; Smilkov et al., 2017). A notable line of work along this direction, called algorithmic recourse (Wachter et al., 2018; Ustun et al., 2019), provides each individual who received an undesirable label (e.g., one whose loan request was denied) with a minimum-cost improvement to achieve the desired label.
An important weakness of much of the work on algorithmic recourse is the assumption that models are fixed and do not change. In practice, models are periodically updated to reflect changes in data or environment, causing a recourse to become invalid, i.e., following it may not yield a desirable outcome (Dominguez-Olmedo et al., 2022). To alleviate this problem, Upadhyay et al. (2021) proposed a framework that is robust to adversarial model changes and provided an algorithm, ROAR, to compute robust recourses. However, guaranteeing robustness in such a setting leads to significantly increased cost (Pawelczyk et al., 2023a).
Our main goal is to design more practical recourse algorithms that balance robustness and cost. To achieve this goal, we move beyond the adversarial setting and leverage the learning-augmented framework (Mitzenmacher and Vassilvitskii, 2020), which has been used in a surge of recent work to overcome the limitations of adversarial (i.e., worst-case) analysis. Specifically, rather than assuming that the designer has no information regarding the model changes, we assume they can at least formulate some predictions regarding these changes. However, these predictions are unreliable, so our goal is to compute recourses that are near-optimal if the predictions are accurate (consistency), while maintaining good performance even in the worst case, i.e., even when the predictions are highly inaccurate (robustness).
Our Results We first adapt the learning-augmented framework and the objectives of consistency and robustness to algorithmic recourse, and we provide a computationally efficient algorithm aiming to optimize these two objectives, as well as their convex combinations. We then formally prove that our algorithm is always optimal for robustness and consistency for any generalized linear model. Note that this is a non-convex problem and, to the best of our knowledge, this is the first optimal algorithm for any robust recourse problem (prior work used algorithms which, as we show, return local rather than global optima). Therefore, this is a contribution to the broader robust recourse literature, beyond the learning-augmented setting. We then experimentally evaluate the algorithm’s performance on a variety of both linear and non-linear models.
@@ 1.1 Related Work
The emerging literature on the interpretability and explainability of machine learning systems mainly advocates for two main approaches. The first approach aims to build inherently simple or interpretable models such as decision lists (Lakkaraju et al., 2016) or generalized additive models (Wang et al., 2017; Yang et al., 2017). These approaches provide global explanations for the deployed models. The second approach attempts to explain the decisions of complex black-box models (such as deep neural networks) only on specific inputs (Ribeiro et al., 2016; Lundberg and Lee, 2017; Smilkov et al., 2017; Sundararajan et al., 2017; Selvaraju et al., 2017; Agarwal et al., 2021; Koh and Liang, 2017; Barshan et al., 2020; Ehyaei et al., 2024). These approaches provide a local explanation of the model and are sometimes referred to as post-hoc explanations.
Recourse is a post-hoc counterfactual explanation that aims to provide the lowest cost modification that changes the prediction for a given input with an undesirable prediction under the current model (Wachter et al., 2018; Ustun et al., 2019; Looveren and Klaise, 2021; Rawal and Lakkaraju, 2020; Slack et al., 2021; Pawelczyk et al., 2023a; Garg et al., 2025). Since its introduction, different formulations have been used to model the optimization problem in recourse (see (Verma et al., 2020) for an overview). Wachter et al. (2018) and Pawelczyk et al. (2023a) considered score-based classifiers and defined modifications to help instances achieve the desired scores. On the other hand, for binary classifiers, Ustun et al. (2019) required the modification to result in the desired label. Roughly speaking, the first setting can be viewed as a relaxation of the second setting, and we follow the second formulation in our work.
The follow-up works on the problem study several other aspects such as focusing on specific models such as linear models (Ustun et al., 2019) or decision-trees (Kanamori et al., 2024; Bewley et al., 2024), understanding the setting and its implicit assumptions and implications (Barocas et al., 2020; Venkatasubramanian and Alfano, 2020; Fokkema et al., 2024; Gao and Lakkaraju, 2023), attainability or actionability (Joshi et al., 2019; Pawelczyk et al., 2020; Karimi et al., 2020a; Ustun et al., 2019), imperfect causal knowledge (Karimi et al., 2020b), fairness in terms of cost of implementation for different subgroups (Guldogan et al., 2023; Gupta et al., 2019; Heidari et al., 2019), repeated dynamics (Fonseca et al., 2023; Bell et al., 2024; Ehyaei et al., 2024), temporal data (Buliga et al., 2025) and providing model-agnostic explanations for recourse by computing responsiveness scores (Cheon et al., 2025). Extending our work to account for these different aspects of recourse is left for future work.
The most closely related work to us is by Upadhyay et al. (2021), which initiated the study of robust recourse and proposes an algorithm called RObust Algorithmic Recourse (ROAR). We study the same framework for our robustness analysis. Nguyen et al. (2022) proposed a new framework for robust recourse that uses a different objective than ROAR and focuses on data shifts instead of model shifts. They call their framework Robust Bayesian Recourse (RBR) and we refer to their algorithm with the same abbreviation. The RoCourseNet algorithm (Guo et al., 2023) also provided robust recourse, though a direct comparison with this algorithm is not possible, as it is an end-to-end approach, i.e., it simultaneously optimizes for the learned model and robust recourse while the initial model is fixed in our approach, just like in (Upadhyay et al., 2021; Nguyen et al., 2022). Similar to the literature on algorithmic recourse, the literature on robust algorithmic recourse has also studied specialized settings (for different model classes, cost functions, and model changes) and additional desiderata such as fairness (Leofante and Potyka, 2024; Yetukuri et al., 2024; Jiang et al., 2024a; Nguyen et al., 2023; Dutta et al., 2022; Mochaourab et al., 2021). Neither approach uses predictions, though we compare our algorithm with both ROAR and RBR when studying the trade-off between validity and cost of robust recourse in the absence of predictions. Very recently, Turbal et al. (2025) provides provably robust algorithms for recourse under predictive multiplicity (Marx et al., 2020) when the set of models corresponds to an ellipsoidal approximation of the Rashomon set (Breiman, 2001) (i.e., models that are similar in performance on the data distribution (Semenova and Rudin, 2019; Xin et al., 2022)). However, in our setting, the set of adversarial models might perform significantly differently on the data distribution. Even more recently Kyaw et al. (2026) introduced algorithms for robust recourse under linear models and $L^{p}$-bounded model changes for the case of $p<\infty$ and showed that this restriction on the adversarial model changes can cause the price of recourse to decrease compared to our adversarial model. See (Jiang et al., 2024b) for a recent survey on robust recourse.
In addition to the above works, Pawelczyk et al. (2023b) studies how model updates due to the “right to be forgotten can affect recourse validity. Dominguez-Olmedo et al. (2022) showed that minimum cost recourse solutions are provably not robust to adversarial perturbations in the model and then presented robust recourse solutions for linear and differentiable models. Black et al. (2022) observe that recourse in deep models can be invalid by small perturbations and suggest that the model’s Lipschitzness at the counterfactual point is the key to preserving validity. Hamman et al. (2023) proposed the notion of “naturally-occurring model change, and provide recourse with theoretical guarantees on the validity of the recourse.
Recourse is also closely related to adversarial training or robust machine learning (Madry et al., 2018; Wong and Kolter, 2018; Athalye et al., 2018). Pawelczyk et al. (2022) studied the connections between various recourse formulations and adversarial training literature. In fact, ROAR (Upadhyay et al., 2021) builds on gradient-based methods originally developed for adversarial training. The convergence of such algorithms has been studied extensively under various assumptions (see, e.g., (Wang et al., 2019)). Theoretical guarantees for adversarial training typically rely on specific data distributions and model classes. For instance, with data drawn from a mixture of Gaussians, the optimal robust classifiers are linear (Li et al., 2020). Moreover, Awasthi et al. (2019) propose an algorithm for learning robust linear classifiers under the realizability assumption and establish hardness results for robust learning of degree-2 polynomial threshold functions. Beyond linear models, theoretical analyses have also been extended to other classes such as decision trees (Vos and Verwer, 2021; 2022). To our knowledge, however, none of these algorithms directly address our setting.
The learning-augmented framework is applied to a wide variety of settings, aiming to provide a refined understanding of the performance guarantees that are achievable beyond the worst case. Unlike approaches in robust optimization that utilize uncertainty sets (Ben-Tal et al., 2009), the learning augmented framework does not impose any assumptions on the quality of the prediction. One of the main application domains is the design of algorithms (e.g., online algorithms (Lykouris and Vassilvitskii, 2018; Purohit et al., 2018)), but it has also been used in data structures design (Kraska et al., 2018), mechanisms interacting with strategic agents (Agrawal et al., 2022; Xu and Lu, 2022), and privacy-preserving methods for processing sensitive data (Khodak et al., 2023). This is already a vast and rapidly growing literature; see (Lindermayr and Megow, 2024) for a frequently updated and organized list of related papers.
@@ 2 Preliminaries
Consider a predictive model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\theta\in\Theta\subseteq\mathbb{R}^{d}$, which maps instances (e.g., loan applicants) from a feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$ to an outcome space $\mathcal{Y}=\{0,1\}$. The values of 0 and 1 represent undesirable and desirable outcomes (e.g., loan denial or approval), respectively. If a model $f_{\theta_{0}}$ yields an undesirable outcome for some instance $x_{0}\in\mathcal{X}$, i.e., $f_{\theta_{0}}(x_{0})=0$, the goal is to suggest an optimal recourse: the least costly way to modify $x_{0}$ (how an applicant should strengthen their application) so that the resulting instance $x^{\prime}$ would achieve the desirable outcome under $f_{\theta_{0}}$, i.e., $f_{\theta_{0}}(x^{\prime})=1$. Given a cost function $c:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ quantifying the cost $c\left(x^{\prime},x_{0}\right)$ required to modify $x_{0}$ to $x^{\prime}$, the recourse is defined as the following optimization problem (Upadhyay et al., 2021):
@@
where $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ is a convex loss function that is decreasing in its first argument (such as binary cross entropy or squared loss) and captures the extent to which the condition $f_{\theta_{0}}(x^{\prime})=1$ is violated and $\lambda\geq 0$ is a regularizer that balances the degree of violation from the desirable outcome and the cost of modifying $x_{0}$ to $x^{\prime}$. The regularizer $\lambda$ can be decreased gradually until the desired outcome is reached. In this work, following the approach of (Upadhyay et al., 2021; Rawal and Lakkaraju, 2020), we assume the cost function is the $L^{1}$ distance, i.e. $c(x,x^{\prime})=\|x-x^{\prime}\|_{1}$.
@@
Equation (1) assumes that the parameters of the model remain the same over time, but in practice, predictive models may be periodically retrained and updated (Upadhyay et al., 2021). These updates can cause a recourse that is valid in the original model (i.e., one that would lead to the desirable outcome in that model) to become invalid in the updated model (Dutta et al., 2022; Black et al., 2022). It is, therefore, natural to require a recourse solution whose validity is robust to (slight) changes in the model parameters.
In line with prior work on robust recourse (Upadhyay et al., 2021; Black et al., 2022), we assume that the parameters of the updated model can be any $\theta^{\prime}\in\Theta_{\alpha}$, where $\Theta_{\alpha}=\{\theta:\|\theta-\theta_{0}\|_{\infty}\leq\alpha\}\subseteq\Theta$ is a “neighborhood” around the parameters, $\theta_{0}$, of the original model, defined using the $L^{\infty}$ distance and a known parameter $\alpha$. Given $\theta_{0}$ and $\Theta_{\alpha}$, the robust solution would be to choose a recourse $x_{r}$ that minimizes the price assuming the parameters of the updated model $\theta^{\prime}\in\Theta_{\alpha}$ are chosen adversarially, i.e.,
@@ 3 Learning-Augmented Framework
Choosing the robust recourse $x_{r}$ according to (3) optimizes the price against an adversarially chosen $\theta^{\prime}\in\Theta_{\alpha}$, but this price may be much higher than the optimal price if we knew $\theta^{\prime}$. This is due to the overly pessimistic assumption that the designer has no information regarding the realized $\theta^{\prime}\in\Theta_{\alpha}$. To overcome similarly pessimistic results in other domains, the learning-augmented framework (Mitzenmacher and Vassilvitskii, 2020) provides a more refined analysis by assuming the designer is equipped with an unreliable prediction and then seeks to achieve near-optimal performance whenever the prediction is accurate while simultaneously maintaining some robustness even if the prediction is arbitrarily inaccurate.
@@
The robustness evaluates the worst-case price of $x^{\prime}$ against an adversarial change of the model and then compares it to the price of $x_{r}$. The robustness is always at least 0 (achieved by $x^{\prime}=x_{r}$), and lower is more desirable. While prior work (Upadhyay et al., 2021) measures robustness in absolute terms, we evaluate it relative to $x_{r}$ to enable a direct comparison between robustness and consistency.
@@ 3.1 Computing Robust/Consistent Recourses
We next provide Algorithm 1 to optimize Objective (7). We introduce a few additional notations. We use $\operatorname{sgn}(s)$ which is 1 when $s>0$, 0 when $s=0$, and 1-1 when $s<0$. When applied to a vector, the $\operatorname{sgn}$ is applied element-wise. For an integer $n\in\mathbb{N}$, $n]:=\{1,\ldots,n\}$.
@@
4:  Initialize worst-case $\theta^{\prime}[i]$ based on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{x^{\prime}}}[i]$ and $\theta[i]$
5:  $\textsc{Active}\leftarrow\textsc{Active}\setminus\{i\}$ if $x^{\prime}_{i}$ cannot improve objective (7)
@@
7:  $i,\Delta\leftarrow$ FindOptimalDimensionAndUpdate($x_{0}$, $\theta^{\prime}$, $\ell$, $c$, $\beta$, $\hat{\theta}$, $x^{\prime}$)
@@
11:   $x^{\prime}[i]\leftarrow x^{\prime}[i]+\Delta$ $\triangleright$ Update $x^{\prime}[i]$
@@
Input: $x_{0}$, $\theta^{\prime}$, $\ell$, $c$, $\beta$, $\hat{\theta}$, and $x^{\prime}$Output: $i,\Delta$
@@
4:  $x^{*}\in\arg\min_{x:x[j]=x^{\prime}[j]\forall j\neq i}K(x)$ $\triangleright$ Minimizer of $K$ if only dimension $i$ could change
5:  if $x^{*}[i]<x^{\prime}[i]\leq x_{0}[i]$ or $x^{*}[i]>x^{\prime}[i]\geq x_{0}[i]$ then
6:   $O[i]\leftarrow K(x^{\prime})-K(x^{*})$ $\triangleright$ Update if change is consistent with prior changes in $x^{\prime}[i]$
@@
Algorithm 1 starts by approximating $f_{\theta}$ locally at $x_{0}$ with a linear model e.g., using LIME (Ribeiro et al., 2016). The algorithm then computes the worst-case model $\theta^{\prime}$ for the “default” recourse of $x^{\prime}=x_{0}$ (lines 3-5). Then, facing $\theta^{\prime}$, the algorithm greedily modifies $x^{\prime}$ while simultaneously updating $\theta^{\prime}$ to ensure that it remains the worst-case model for $x^{\prime}$ (lines 6-17). More specifically, in each iteration of the while loop, for the current $\theta^{\prime}$, the algorithm first calls the subroutine FindOptimalDimensionAndUpdate (line 1), implemented in Algorithm 2. This subroutine searches over each dimension and returns the dimension $i$, and the change $\Delta$ in that dimension that minimizes the objective (7), assuming $\theta^{\prime}$ and all other coordinates of $x^{\prime}$ are fixed. Suppose the $\Delta$ returned by Algorithm 2 is 0, meaning no improvement can be made on any dimension, then Algorithm 1 terminates (line 9). If $|\Delta|>0$, but applying this change to $x^{\prime}[i]$ does not cause $x^{\prime}[i]$ to change sign (i.e., $x^{\prime}[i]+\Delta$ has the same sign as $x^{\prime}[i]$), then the update is applied; in that case, the adversarial model $\theta_{0}$ does not need to be updated (line 11). On the other hand, if the recommended change flips the sign of $x^{\prime}[i]$, this would cause the adversarial response to change as well. The algorithm instead applies this change all the way up to $x^{\prime}[i]=0$ (i.e., up until the sign change), which maintains the same adversarial response. Then, $\theta^{\prime}$ is updated with $\theta^{\prime}[i]\leftarrow\theta[i]+\alpha\cdot\operatorname{sgn}(x_{0}[i])$ (lines 13-15), so that any subsequent changes of $x^{\prime}[i]$ beyond $x^{\prime}[i]=0$ are calculated with respect to the correct adversarial response. If, this update of $\theta^{\prime}$ causes the sign of $\theta^{\prime}[i]$ to change (i.e., if $|\theta[i]|>\alpha$), then no further change in this dimension can yield an improvement on our objective and $i$ is removed from the Active set (line 17).
Algorithm 1 is efficient since (1) the linear approximation of $f_{\theta_{0}}$ can be computed in polynomial time (Ribeiro et al., 2016), and (2) the while loop runs for $O(d)$ iterations and the running time of each iteration is dominated by the runtime of subroutine FindOptimalDimensionAndUpdate. This subroutine runs in polynomial time (e.g., by running at most $d$ one-dimensional gradient descents (Boyd and Vandenberghe, 2014)). See Appendix B for more details, and Appendix C.7 for the scalability experiments.
@@
Proof Sketch.
We provide a sketch of the proof for the case of $\beta=1$ (i.e., for optimal robustness), and defer the full proof and the analysis of the case of $\beta=0$ (i.e., for optimal consistency) to Appendix B.2.
We start with Observation B.2, which points out that we can assume $J(x^{\prime},\theta^{\prime})$ is a function only of the cost $\|x^{\prime}-x_{0}\|_{1}$ of $x^{\prime}$ and its inner product with its worst-case (adversarial) model, $x^{\prime}\cdot\theta^{\prime}$. Specifically, it is a linear increasing function of the former and a convex decreasing function of the latter. Using this observation, we view the problem as choosing a recourse $x^{\prime}$ and suffering a cost of $\|x^{\prime}-x_{0}\|_{1}$, aiming to maximize $x^{\prime}\cdot\theta^{\prime}$, and an adversary then chooses $\theta^{\prime}$ aiming to minimize this inner product. The actual form of $J(\cdot)$ determines the extent to which we may want to trade off the cost of recourse $x^{\prime}$ for an increase in the inner product, but the convexity with respect to the inner product suggests a decreasing marginal gain with respect to the latter.
To better understand the structure of the adversarial response, Lemma B.3 shows that for any given recourse $x$, the adversarial model $\theta^{\prime}$ is essentially equal to $\theta_{0}-\alpha\cdot\operatorname{sgn}(x)$. In other words, the adversary shifts each coordinate $i$ by $\alpha$ (the maximum shift that is allowed while remaining within $\Theta_{\alpha}$), and the direction is determined by whether $x[i]$ is positive or negative. As a result, we can assume that the adversarial choice of $\theta^{\prime}$ for each dimension $i$ is always either $\theta_{0}+\alpha$ or $\theta_{0}-\alpha$. This also allows us to partition $\mathcal{X}$, the space of all possible recourses, into regions that would face the same adversary. Specifically, if $x$ and $x^{\prime}$ are such that $\operatorname{sgn}(x)=\operatorname{sgn}(x^{\prime})$, then their worst-case model is the same.
Our algorithm starts from $x_{0}$ and gradually changes it until it reaches the robust recourse $x_{r}$. To determine its first step, the algorithm computes the worst-case model $\theta^{\prime}$ for $x_{0}$ using the formula provided by Lemma B.3. If we were to assume that the adversary would remain fixed at $\theta^{\prime}$ no matter how we change the recourse, then computing the optimal recourse would be easy: we would identify the dimension $i$ with the largest $|\theta^{\prime}[i]|$ value, and we would change only this dimension of $x_{0}$ (This would in turn simplify how $i$ in line 1 can be computed for the specific case of $\beta=1$. See Appendix B.2). This is optimal because changing some other dimension $j$ by $\Delta$ would increase our cost by $\Delta$ and increase the inner product by $\Delta\cdot|\theta^{\prime}[j]|$ against a fixed adversary. How much we should change that dimension would then be determined by solving the optimization problem in line 1. However, if this change “flipped” the sign of $x^{\prime}[i]$, this would also change the adversary, potentially compromising the optimality of $x^{\prime}$.
Our proof shows that a globally optimal recourse can be computed by myopically optimizing $x^{\prime}$ until the adversary needs to be updated, and then repeating the same process until the marginal gain in the inner product is outweighed by the marginal increase in cost. This is partly due to the fact (shown in Lemma B.5) that the order in which this myopic approach considers the dimension of the recourse to change is optimal, exhibiting a decreasing sequence of $|\theta^{\prime}[i]|$ values. ∎
Note that $\beta=1$ corresponds to computing the optimal robust recourse. Even for generalized linear models, the objective function in Equation (7) is non-convex when $\beta=1$ (see Appendix B.1). Hence, prior gradient-based approaches (Upadhyay et al., 2021; Nguyen et al., 2022) cannot guarantee optimality. To the best of our knowledge, Algorithm 1 is the first optimal algorithm for any robust recourse formulation. This optimality guarantee relies on the assumption that the single-dimensional convex problem in line 4 of Algorithm 2 can be solved in polynomial time. This can be done, for example, when the loss function $\ell$ is known by deriving first-order conditions and verifying optimality conditions. In general, gradient-based approaches can be used to converge to an optimal solution, in which case the solution provided by Algorithm 1 will also converge to the optimal robust or consistent recourse.
The idea of approximating a non-linear function locally has also been used in prior work (Upadhyay et al., 2021; Rawal and Lakkaraju, 2020). In Section 4, we empirically compare the performance of our algorithm to prior work in linear and non-linear settings. To handle feasibility constraints or categorical features, the recourse of Algorithm 1 can be post-processed (e.g., by projection) to guarantee feasibility (Upadhyay et al., 2021; Nguyen et al., 2022; Guo et al., 2023). See Appendix C.4.